How do I find the base of a triangle? The area of the triangle is 200 ft. squared, and the height is 12ft.
Answer by checkley71(8403)

AREA=B*H/2
200=B*12/2
200=6B
B=200/6
B=33 1/3 IS THE BASE
PROOF
200=(33 1/3)12/2
200=(33*12+12/3)/2
200=(396+4)/2
200=400/2
200=200